2. Cell death due to hyperthermia has been described by first-order kinetics with an Arrhenius description of k(T). That is ITable 1: Table for Problem # 2. Cell death half-times for heating at various temperatures Temperature Time for half of the ceSolve for B and C
Answer #1

solution? - frest-order expression . en (A ) = -kt (6) All Joe siest ordeę reaction I tip = 0.693 so from here we can says- STaking Ti= 42°C = 315k T BYSOC = 318 K (Ain), = 31000see (1/2)2 = 17000 see- Put all values! - esp (elety lata-sha)) - ( Ca =New! em Condom --K220+ per = erp (- coussioʻx 40) = 0.997 NO